OpenStudy (anonymous):

Solve 3t(t+2)-3t^2 for t=19 I got 836.

OpenStudy (campbell_st):

so its \[3 \times 19(19 + 2) - 3 \times 19^2 = 57 \times 21 - 3 \times 361\] I think you need to check your answer

OpenStudy (dangerousjesse):

Yeah, I got a way smaller answer. Just work out the problem by swapping t with 19 and using PEMDAS.

OpenStudy (anonymous):

I forgot about PEMDAS so I redid the problem and got 114.

OpenStudy (dangerousjesse):

Yes, that's the answer :)

OpenStudy (anonymous):

thank you! :)

OpenStudy (dangerousjesse):

No problem.